OpenStudy (abbycross167):

Find the point-slope form of the equation that goes through these two points Points ~ (1,20) (8,4.5) Equation ~ -31/14x+311/14 Can someone please help me?

OpenStudy (addicted2anime):

ok so for the two sets of numbers the equation is m=\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (addicted2anime):

plug the set of numbers in

OpenStudy (addicted2anime):

y1-y2= 20-4.5 =15.5 x1-x2= 1-8=-7

OpenStudy (addicted2anime):

wait don't listen 2 me, confused it with slope

OpenStudy (abbycross167):

haha umm ok yes sir/ma'am

OpenStudy (abbycross167):

thank you for your help tho

OpenStudy (addicted2anime):

np i'm nuala and im a girl lol